OpenStudy (anonymous):

Simplify the given expression. 3 over 4 i negative 3 over 4 3 I over 4 negative 3 I over 4 12 I over 16 I squared

OpenStudy (anonymous):



OpenStudy (anonymous):

multiply top and bottom by either \(i\) or \(-i\) your choice

OpenStudy (anonymous):

\[\frac{3}{4i}\times \frac{i}{i}=\frac{3i}{-4}=-\frac{3}{4}i\]

OpenStudy (anonymous):

so its